How many liters of 10% alcohol solution and 5% alcohol solution must be mixed to obtain 40 liters of 8% alcohol solution?

Answer by lwsshak3(11628)

How many liters of 10% alcohol solution and 5% alcohol solution must be mixed to obtain 40 liters of 8% alcohol solution?
***
let x=amt of 10% solution to mix
40-x=amt of 5% solution to mix
..
10%x+5%(40-x)=8%*40
.10x++2-.05x=3.2
.05x=1.2
x=24
40-x=16
amt of 10% solution to mix=24 liters
amt of 5% solution to mix=16 liters